OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x − 1) −12? x −5 x −5 −5 x A number line with a closed circle on negative 5 and shading to the right.

OpenStudy (anonymous):

what is the inequality sign....you need to re- write your inequality .

OpenStudy (anonymous):

less than or equal to.

OpenStudy (anonymous):

Is this the inequality.... 2(x - 1) <= -12 ?

OpenStudy (anonymous):

No, it's less than or equal to, not greater than.

OpenStudy (anonymous):

that is the less then sign

OpenStudy (anonymous):

Oh, my bad greater than. I was looking at the wrong one, lol, sorry. It is greater than or equal to.

OpenStudy (anonymous):

2(x - 1) <= - 12 2x - 2 <= - 12 2x <= -12 + 2 2x <= -10 x <= -5 your answer choices do not contain inequality signs..... the last one is not the correct answer (I know that for sure)